# Temperature profiles of accretion disks in luminous active galactic nuclei derived from ultraviolet spectroscopic variability

### v1

@@ 6 Conclusion
We construct ultraviolet ensemble SFs of luminous AGNs grouped by luminosity and continuum wavelength using spectrophotometric data obtained through the SDSS-RM project. Assuming that all AGNs share a single universal SF after $\Delta t$ is normalized by the characteristic timescale ($\tau$), we fit all the ensemble SFs to investigate how the characteristic timescale depends on luminosity and wavelength in the form $\tau\propto L^{a}\lambda_{\rm RF}^{b}$. Our main findings are.
@@
  • The dependence of $\tau$ on continuum wavelength ($b=1.42\pm 0.09$) departs by more than $6\,\sigma$ from the standard disk model ($b=2$), revealing a shallower color (steeper temperature) profile.
  • Highly accreting AGNs and Balmer continuum have a negligible effect on measurements of $b$, strongly suggesting that the discrepancy between the observed and predicted value of $b$ is genuine.
  • Interestingly, our estimation of $b$ based on spectroscopic variability is in broad agreement with previous continuum RM and microlensing analysis based on spectroscopic monitoring data. This suggests that spectroscopic data may yield more robust results on the radial temperature or color profiles than photometric data.
@@ Acknowledgements.
Table 1: Properties for the Subsamples
@@
666 Col. (1): Central wavelength of continuum windows in the rest frame. Col. (2): Group divided by bolometric luminosity: 0 = $45.4<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 45.65$, 1 = $45.65<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 45.9$, 2 = $45.9<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 46.15$, 3 = $46.15<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 46.4$, 4 = $46.4<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 46.65$, 5 = $46.65<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 46.9$, and 6 = $46.9<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 47.15$. Col. (3): Number of objects; in parentheses is the number of objects with BH mass estimates. Col. (4): Average redshift. Col. (5): Average bolometric luminosity. Col. (6): Average BH mass. Col. (7): Average Eddington ratio.
Table 2: Best-fit parameters for $\tau\propto L^{a}\,\lambda^{b}_{\rm RF}$
Subsample $a$ $b$
(1) (2) (3)
All $0.50\pm 0.03$ $1.42\pm 0.09$
$0.01<\lambda_{\rm Edd}\leq 0.3$ $0.46\pm 0.07$ $1.32\pm 0.15$
All with $\lambda_{\rm RF}<3000\,{\rm\AA}$ $0.50\pm 0.04$ $1.45\pm 0.18$

### v2

@@ 6 Conclusion
We constructed ultraviolet ensemble SFs of luminous AGNs grouped by luminosity and continuum wavelength using spectrophotometric data obtained through the SDSS-RM project. Assuming that all AGNs share a single universal SF after $\Delta t$ is normalized by the characteristic timescale ($\tau$), we fit all the ensemble SFs to investigate how the characteristic timescale depends on luminosity and wavelength in the form $\tau\propto L^{a}\lambda_{\rm RF}^{b}$. Our main findings are:
@@
  • The dependence of $\tau$ on the continuum wavelength ($b=1.42\pm 0.09$) departs by more than $6\,\sigma$ from the standard disk model ($b=2$), revealing a shallower color (steeper temperature) profile.
  • Highly accreting AGNs and the Balmer continuum have a negligible effect on measurements of $b$, strongly suggesting that the discrepancy between the observed and predicted value of $b$ is genuine.
  • Interestingly, our estimation of $b$ based on spectroscopic variability is in broad agreement with previous continuum RM and microlensing analyses based on spectroscopic monitoring data. This suggests that spectroscopic data may yield more robust results on the radial temperature or color profiles than photometric data.
@@ Acknowledgements.
Appendix A AGN properties of the sample
The properties of AGNs and the number of objects for each subsample are listed in Table A.1.
Table 2: Properties of the subsamples.
@@ Appendix A AGN properties of the sample
666 Col. (1): Central wavelength of continuum windows in the rest-frame. Col. (2): Group divided by bolometric luminosity: 0 = $45.4<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 45.65$, 1 = $45.65<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 45.9$, 2 = $45.9<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 46.15$, 3 = $46.15<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 46.4$, 4 = $46.4<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 46.65$, 5 = $46.65<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 46.9$, and 6 = $46.9<$ $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 47.15$. Col. (3): Number of objects; in parentheses is the number of objects with BH mass estimates. Col. (4): Average redshift. Col. (5): Average bolometric luminosity. Col. (6): Average BH mass. Col. (7): Average Eddington ratio.